OpenStudy (anonymous):

y=2x^2-7x+7 the function above crosses the x-axis once. True or false?

OpenStudy (bahrom7893):

false

myininaya (myininaya):

y=2x^2-7x+7 2x^2-7x+7=0

OpenStudy (anonymous):

Thank you:)

OpenStudy (bahrom7893):

as I showed before it doesn't cross the x axis at all..

myininaya (myininaya):

solve that equation

OpenStudy (anonymous):

Isn't this the same question?

OpenStudy (bahrom7893):

http://www.wolframalpha.com/input/?i=Graph%28y%3D2x%5E2-7x%2B7%29 it is..

myininaya (myininaya):

\[x=\frac{7 \pm \sqrt{49-4(2)(7)}}{4}\] since this is imaginary there are no x-intercepts the graph does not touch x-axis

OpenStudy (anonymous):

Yes, but how can we see that mathematically? A function crosses the x axis, when y=0. So, let's set the equation equal to 0 and solve for x. That will tell us what places the graph crosses. So we get 2x^2-7x+7 = 0 Using the quadratic formula, we get x = (7 +/- sqrt(49 - 56)) /4 = (7 +/- sqrt(-7))/4 So since the sqrt is of a negative number, there is no real solution.